# MARVEL: Unlocking the Multi-Modal Capability of Dense Retrieval via Visual Module Plugin

### v3

@@ A.5 Additional Ablation Studies on MARVEL
As shown in Table 11, the relevance modeling between queries and image documents heavily depends on the image caption, which is also observed in previous work Liu et al. (2023b). The image features contribute to approximately 1% improvements in the image retrieval task, demonstrating the effectiveness of image features in helping model better understand the image documents.
@@ A.6 Case Studies
For the first case, UniVL-DR conducts shallow keyword matching and returns text documents that are related to “animal” and “Peace” mentioned in the query, which are unrelated to the query. MARVEL can better understand that “Peace and Plenty” is a famous painting and retrieve more related images and text documents for users. In the second case, UniVL-DR, MARVEL-DPR, and MARVEL-ANCE all return images or text documents related to “promotion ideas”. Notable, MARVEL can better understand the user’s question and return the modal that the user expects. MARVEL-ANCE introduces a variety of sales promotion strategies rather than matching on “promotion” keywords. It shows the effectiveness of MARVEL in better fusing the retrieval results from different modalities, which thrives on universal multi-model document encoding.

### v4

@@ A.5 Additional Ablation Studies on MARVEL
As shown in Table 10, the relevance modeling between queries and image documents heavily depends on the image caption, which is also observed in previous work Liu et al. (2023b). The image features contribute to approximately 1% improvements in the image retrieval task, demonstrating the effectiveness of image features in helping the model better understand the image documents.
A.6 Learned Semantics of Image Features
As shown in Table 11, replacing encoded image features with $k$-NN token embeddings generally outperforms the retrieval model using randomly selected token embeddings. It demonstrates that the visual plugin module effectively maps image semantics in the input space of the language model, and the ability to keep growing with more token embeddings (5-NN). However, the retrieval performance significantly drops when employing $k$-NN token embeddings to replace the image features, compared to the MARVEL model. It demonstrates the role of encoded image features beyond the semantic representations of word embeddings. The encoded image features may act as a kind of prompt, encouraging language models to capture image semantics Merullo et al. (2023).
A.7 Retrieval Efficiency of MARVEL
In this section, we compare the retrieval efficiency of MARVEL with other baselines on the same device, as shown in Table 12.
MARVEL follows the general dense retrieval framework Karpukhin et al. (2020) for efficient document retrieval. It encodes the entire corpus offline and constructs the document index using FAISS Johnson et al. (2019) for online searches. While the offline encoding time for queries and image/text documents in MARVEL is longer than that in UniVL and other baseline models, this encoding process does not impact retrieval efficiency.
When comparing the retrieval latency of these models in retrieving the top 100 relevant documents for each query, MARVEL’s retrieval time is comparable to its base model, T5-ANCE, which has a retrieval time of 48.3 ms. Furthermore, the retrieval time for document encoding using the CLIP model is less than that of models such as T5 and BERT. This demonstrates that the retrieval time is influenced by the dimensionality of the embeddings, and MARVEL’s architecture does not introduce any additional retrieval latency.
@@ A.8 Case Studies
For the first case, UniVL-DR conducts shallow keyword matching and returns text documents that are related to “animal” and “Peace” mentioned in the query, which are unrelated to the query. In contrast, MARVEL can better understand that “Peace and Plenty” is a famous painting and retrieve more related images and text documents for users. In the second case, UniVL-DR, MARVEL-DPR, and MARVEL-ANCE all return documents related to “promotion ideas”. Notable, MARVEL can better understand the user’s query and return the expected modality. MARVEL-ANCE introduces a variety of sales promotion strategies rather than matching on “promotion” keywords. It shows the effectiveness of MARVEL in better fusing the retrieval results from different modalities, which thrives on universal multi-modal document encoding.